### v2

@@ 7 Adversarial Setting
Suppose for contradiction that $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})\neq\emptyset$ and $N_{{\mathcal{A}}}(r\mid S^{\prime})\not\subseteq N_{{\mathcal{A}}}(r\mid S)$, and let $s\in N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\setminus N_{{\mathcal{A}}}(r\mid S)$. Since $r$ is the first arrived request, all servers in $S$ (resp. $S^{\prime}$) are available upon the arrival of $r$, and hence, $N_{{\mathcal{A}}}(r\mid S)$ (resp. $N_{{\mathcal{A}}}(r\mid S^{\prime})$) contains all servers in $S$ (resp. $S^{\prime}$) closest to $r$. Moreover, note that $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq S^{\prime}\subseteq S$. On one hand, since $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, $\delta(s,r)=\delta(s^{\prime},r)$. On the other hand, since $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\notin N_{{\mathcal{A}}}(r\mid S)$, $\delta(s,r)<\delta(s^{\prime},r)$, leading to a contradiction. ∎
We analyze the expected cost of ${\mathcal{A}}$ under both cases in Lemma 11 separately. We start with the easier case where $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$, which implies that $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$ and $\delta(r,s)<\delta(r,s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. By $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$, we know that $M_{{\mathcal{A}}}(r\mid S)\notin S^{\prime}$ and hence $S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\}\subseteq S^{\prime}\subseteq S\setminus\{M_{{\mathcal{A}}}(r\mid S)\}$. By the inductive hypothesis, $\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})$. As a result,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)$ $=\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S))+\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})]$
$<\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S^{\prime}))+\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})]$
$=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\,,$
concluding the proof.
@@
$P_{S}(s)$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S)=s\right\},$ $Q_{S}(s)$ $=\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{s\},R\setminus\{r\})\,,$
$P_{S^{\prime}}(s^{\prime})$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S^{\prime})=s^{\prime}\right\},$ $Q_{S^{\prime}}(s^{\prime})$ $=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{s^{\prime}\},R\setminus\{r\})\,.$
@@
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$
@@
where the second equality holds since $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$ and the inequality holds since $Q_{S}(s)\leq Q_{S^{\prime}}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$ by the inductive hypothesis. Note that for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, we have $S^{\prime}\setminus\{s^{\prime}\}\subseteq S^{\prime}\subseteq S\setminus\{s\}$ and hence $Q_{S}(s)\leq Q_{S^{\prime}}(s^{\prime})$ by the inductive hypothesis. Let $B\triangleq\min_{s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s^{\prime})$, and it follows that $Q_{S}(s)\leq B\leq Q_{S^{\prime}}(s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. Therefore,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$ $\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
@@
where the second inequality holds because $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. This concludes the proof. ∎
7.1 Applications on Hierarchically Separated Tree Metrics
Now, we apply Theorem 9 to algorithms of prior work. Recall that for balanced markets with adversarial servers and requests, the algorithm of [MNP06], abbreviated as the MNP algorithm, is $O(\log^{3}n)$-competitive for general metrics, and the Random-Subtree algorithm of [GL12] is $O(\log n)$-competitive for doubling metrics. We show that the competitive ratios of both algorithms are retained in unbalanced markets.
Both the MNP and Random Subtree algorithms are designed based on the same framework. They first randomly embed the original metric into a Hierarchically Separated Tree (HST) metric, which preserves the distance between every pair of nodes up to an $O(\log n)$ distortion. Next, they show that a randomized greedy subroutine with a specific tie-breaking rule admits a competitive ratio of $\beta$ for the HST metric. Overall, this leads to an $O(\beta\log n)$-competitive algorithm for the original metric. Hence, for our purpose, it suffices to show that the tie-breaking rule of the greedy subroutine for the HST metric is local and non-increasing.
Recall that an HST is a rooted tree satisfying the following property: For each subtree, the distances from the subtree’s root to all leaf nodes in the subtree are identical. Moreover, the metric embedding phase of the algorithms will map each location in the original metric to a leaf node of the HST. From now on, we will restrict our attention to the HST metric with all servers and requests locating at the HST’s leaf nodes.
We start by analyzing the greedy subroutine adopted by the MNP algorithm, which works as follows. Upon the arrival of a new request $r$, assuming the set of initial servers to be $S$, denote $A(r\mid S)$ as the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, and the subroutine matches $r$ to a uniformly random server among all available servers in the subtree rooted at $A(r\mid S)$. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the MNP algorithm satisfies the conditions in Theorem 9.
@@ 7.1 Applications on Hierarchically Separated Tree Metrics
The greedy subroutine in the MNP algorithm admits a local and non-increasing tie-breaking rule.
The proof of Lemma 12 is deferred to Section B.1. The following corollary is a direct consequence of combining Theorem 9 and Lemma 12 with the competitive ratio guarantee of the MNP algorithm in balanced markets [MNP06].
@@
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For an arbitrary metric, an $O(\log^{3}n)$-competitive algorithm exists.
We remark here that for general metrics with adversarial servers and requests, a randomized algorithm with an improved competitive ratio of $O(\log^{2}n)$ exists [BBGN14]. However, this algorithm does not belong to the family of greedy algorithms, and hence, Theorem 9 does not apply.
Next, we analyze the greedy subroutine adopted by the Random-Subtree algorithm, which works as follows. Upon the arrival of a new request $r$, recall that $A(r\mid S)$ denotes the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, assuming the set of initial servers to be $S$. Starting off at $A(r\mid S)$, the subroutine repeatedly moves us to a uniformly random subtree among all subtrees of the current node containing at least one available server until we reach a leaf node, and $r$ is matched to a uniformly random available server at the leaf node. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the Random Subtree algorithm satisfies the conditions in Theorem 9.
@@
The greedy subroutine in the Random Subtree algorithm admits a local and non-increasing tie-breaking rule.
The proof of Lemma 13 is deferred to Section B.2. Combining Theorem 9 and Lemma 13 with the competitive ratio guarantee of the Random Subtree algorithm in balanced markets [GL12], we derive the competitive ratio guarantee of the Random Subtree algorithm in unbalanced markets in the following corollary.
@@
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For doubling metrics, an $O(\log n)$-competitive algorithm exists.
@@ 8 Discussion and Future Directions
In this paper, we present a generic approach for online metric matching to utilize the provided distributional knowledge or predictions of request locations, leading to new results in both settings.
@@
When requests are stochastic, for balanced markets, we achieve an $O(1)$ competitive ratio for the Euclidean space with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for achieving this goal comprise (1) improving the analysis of the algorithm of [GGPW19], which is proved to be $O((\log\log\log n)^{2})$-competitive without any super-constant lower bound known, and (2) applying Theorem 2 and the SOAR algorithm to the simplified model given by [GGPW19]. The main technical barrier underlying the latter approach lies in the difficulty of characterizing $\mathrm{OPT}(n)$ for general metrics and distributions. In Appendix D, we provide further evidence that supports the plausibility of this approach by applying it to achieve a constant competitive ratio with tree metrics and general distributions, recovering a result achieved by the algorithm of [GGPW19].
@@
We also derive an algorithm in unbalanced markets with improved competitive ratios when requests are drawn from smooth distributions over the Euclidean space. Our algorithm admits an optimal constant competitive ratio for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable one with the presence of excess servers, rendering its resolution more appealing.
@@
Moreover, for adversarial servers and requests without predictions, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all the existing algorithms. In this most extensively studied setting, our understanding of the achievable guarantees in unbalanced markets remains very limited, and nothing beyond our results is known. An exciting future direction is to generalize the performance of the existing algorithms with provable guarantees in balanced markets to unbalanced markets.
@@
Several open questions arise from the modeling assumptions in this paper. First, our results assume advance knowledge of $n$, raising the question of whether competitive guarantees are achievable without this information. Next, while we define unbalanced markets as $n\leq m\leq O(n)$ for our competitive ratio results in the stochastic setting, a definition adopted by all prior work, it remains unclear whether these results extend to the cases where $m$ is polynomially or exponentially larger than $n$ [^5] Notice that our results for regret analysis and those in the adversarial setting hold regardless of the number of excess servers. In particular, as $m$ increases, the competitive ratio of our algorithm also increases correspondingly with a context-dependent increasing rate, and when $m=\omega(n)$, our algorithm no longer achieves a constant competitive ratio for the Euclidean space with $d\geq 3$. Moreover, relaxing our assumptions that servers outnumber requests and all requests must be served leads to an interesting variant where request rejection is permitted—a setting that bears similarities to the multi-unit prophet inequality problem [JMZ22]. Finally, it is appealing to relax the assumption that the cost function satisfies the triangle inequality, and we establish positive results for cost functions only approximately satisfying the triangle inequality in Appendix C.

### v3

@@ 7 Adversarial Setting
Suppose for contradiction that $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})\neq\emptyset$ and $N_{{\mathcal{A}}}(r\mid S^{\prime})\not\subseteq N_{{\mathcal{A}}}(r\mid S)$, and let $s\in N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\setminus N_{{\mathcal{A}}}(r\mid S)$. Since $r$ is the first arrived request, all servers in $S$ (resp. $S^{\prime}$) are available upon the arrival of $r$, and hence $N_{{\mathcal{A}}}(r\mid S)$ (resp. $N_{{\mathcal{A}}}(r\mid S^{\prime})$) contains all servers in $S$ (resp. $S^{\prime}$) closest to $r$. Moreover, note that $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq S^{\prime}\subseteq S$. On one hand, since $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, $\delta(s,r)=\delta(s^{\prime},r)$. On the other hand, since $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\notin N_{{\mathcal{A}}}(r\mid S)$, $\delta(s,r)<\delta(s^{\prime},r)$, leading to a contradiction. ∎
We analyze the expected cost of $\mathcal{A}$} under both cases in Lemma 11 separately. We start with the easier case where $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$, which implies that $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$ and $\delta(r,s)<\delta(r,s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. By $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$, we know that $M_{{\mathcal{A}}}(r\mid S)\notin S^{\prime}$ and hence $S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\}\subseteq S^{\prime}\subseteq S\setminus\{M_{{\mathcal{A}}}(r\mid S)\}$. By the inductive hypothesis, $\mathrm{cost}({\mathcal{A}};S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})\leq\mathrm{cost}({\mathcal{A}};S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})$. As a result,
$\mathrm{cost}({\mathcal{A}};S,R)$ $=\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S))+\mathrm{cost}({\mathcal{A}};S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})]$
$<\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S^{\prime}))+\mathrm{cost}({\mathcal{A}};S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})]$
$=\mathrm{cost}({\mathcal{A}};S^{\prime},R)\,$,
as desired.
@@
$P_{S}(s)$ $:=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S)=s\right\},$ $Q_{S}(s)$ $:=\mathrm{cost}({\mathcal{A}};S\setminus\{s\},R\setminus\{r\})\,$,
$P_{S^{\prime}}(s^{\prime})$ $:=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S^{\prime})=s^{\prime}\right\},$ $Q_{S^{\prime}}(s^{\prime})$ $:=\mathrm{cost}({\mathcal{A}};S^{\prime}\setminus\{s^{\prime}\},R\setminus\{r\})\,.$
@@
$\mathrm{cost}({\mathcal{A}};S,R)-\mathrm{cost}({\mathcal{A}};S^{\prime},R)$
@@
where the second equality holds since $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$, and the inequality holds since $Q_{S}(s)\leq Q_{S^{\prime}}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$ by the inductive hypothesis. Note that for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, we have $S^{\prime}\setminus\{s^{\prime}\}\subseteq S^{\prime}\subseteq S\setminus\{s\}$ and hence $Q_{S}(s)\leq Q_{S^{\prime}}(s^{\prime})$ by the inductive hypothesis. Let $B:=\min_{s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s^{\prime})$, and it follows that $Q_{S}(s)\leq B\leq Q_{S^{\prime}}(s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. Therefore,
$\mathrm{cost}({\mathcal{A}};S,R)-\mathrm{cost}({\mathcal{A}};S^{\prime},R)$
$\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
@@
where the second inequality holds since $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. This concludes the proof. ∎
7.1 Applications to Existing Algorithms
Now, we apply Theorem 9 to algorithms of prior work. Recall that for balanced markets and the adversarial setting, the algorithm of [MNP06], abbreviated as the MNP algorithm, is $O(\log^{3}n)$-competitive for general metrics, and the Random-Subtree algorithm of [GL12] is $O(\log n)$-competitive for doubling metrics. We show that the competitive ratios of both algorithms are retained in unbalanced markets.
Both the MNP and Random Subtree algorithms are designed based on the same framework. They first randomly embed the original metric into a Hierarchically well-Separated Tree (HST) metric, which preserves the distance between every pair of locations up to an $O(\log n)$ distortion. Next, they show that a randomized greedy subroutine with a specific tie-breaking rule is $\beta$-competitive for the HST metric. Overall, this leads to an $O(\beta\log n)$-competitive algorithm for the original metric. For our purpose, it suffices to show that the tie-breaking rule of the greedy subroutine for the HST metric is local and non-increasing.
Recall that an HST is an edge-weighted rooted tree satisfying the following property: For each node $x$ of the HST, the distances from $x$ to all leaf nodes in the subtree rooted at $x$ are identical. [^4] The definition of HSTs also includes other conditions, but they are not needed for us. Moreover, the metric embedding phase of the algorithms will map each location in the original metric to a leaf node of the HST. From now on, we will restrict our attention to the HST metric with all servers and requests located at the HST’s leaf nodes.
We start by analyzing the greedy subroutine in the MNP algorithm, which works as follows. Upon the arrival of a new request $r$, assuming the set of initial servers to be $S$, denote $A(r\mid S)$ as the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, and the subroutine matches $r$ to a uniformly random available server in the subtree rooted at $A(r\mid S)$. In the following lemma, we show that this tie-breaking rule is local and non-increasing.
@@ 7.1 Applications to Existing Algorithms
The tie-breaking rule of the greedy subroutine in the MNP algorithm is local and non-increasing.
The proof of Lemma 12 is deferred to Section B.1. The following corollary is a direct consequence of combining Theorem 9 and Lemma 12 with the competitive ratio of the MNP algorithm in balanced markets [MNP06].
@@
In the adversarial setting and unbalanced markets, there exists an $O(\log^{3}n)$-competitive algorithm for general metrics.
We remark here that in the adversarial setting and unbalanced markets, an $O(\log^{2}n)$-competitive algorithm for general metrics is known [BBGN14]. However, the subroutine of this algorithm is not greedy, and hence Theorem 9 does not apply.
Next, we analyze the greedy subroutine in the Random-Subtree algorithm, which works as follows. Upon the arrival of a new request $r$, starting off at $A(r\mid S)$, the subroutine repeatedly moves to a uniformly random subtree of the current node that contains at least one available server, until a leaf node is reached. Then, $r$ is matched to a uniformly random available server at this leaf node. We show in the following lemma that this tie-breaking rule is local and non-increasing.
@@
The tie-breaking rule of the greedy subroutine in the Random Subtree algorithm is local and non-increasing.
The proof of Lemma 13 is deferred to Section B.2. Combining Theorem 9 and Lemma 13 with the competitive ratio of the Random Subtree algorithm in balanced markets [GL12], we obtain the following corollary.
@@
In the adversarial setting and unbalanced markets, there exists an $O(\log n)$-competitive algorithm for doubling metrics.
@@ 8 Discussion and Future Directions
In this paper, we present a generic framework for online metric matching to leverage the provided prediction of request locations, and we apply this framework to achieve new results in the semi-stochastic and prediction settings.
@@
For the semi-stochastic setting and balanced markets, we give an $O(1)$-competitive algorithm for the Euclidean metric with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for accomplishing this goal comprise (1) improving the analysis of the $O((\log\log\log n)^{2})$-competitive algorithm by [GGPW19], or (2) applying Theorem 2 and the SOAR algorithm to the simplified model given by [GGPW19], where $\mathbb{D}$} is uniformly distributed over server locations. The main technical barrier underlying the latter approach lies in characterizing $\mathrm{OPT}_{\mathbb{D}}(n)$ for general metrics. In Appendix C, we further substantiate the potential of this approach by showing that it gives an $O(1)$-competitive algorithm for tree metrics and general distributions, recovering the result by [GGPW19] via the first approach.
@@
Besides, for the semi-stochastic setting and unbalanced markets, we derive an algorithm with improved competitive ratios for the Euclidean metric with smooth distributions. The resulting competitive ratio is optimal for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable with the presence of excess servers, rendering its resolution more appealing.
@@
Moreover, for the adversarial setting without predictions, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all existing algorithms. A natural subsequent step would be to generalize the performance of other algorithms, such as the algorithms by [BBGN14, Rag16], to unbalanced markets.
@@
Several research avenues arise from the modeling assumptions in this paper. First, our results assume advance knowledge of $n$, raising the question of whether competitive guarantees are achievable without this information. Second, the competitive ratios achieved by our algorithm in the semi-stochastic setting for unbalanced markets (Theorem 6) depend polynomially on $\kappa:=m/n$, and hence no longer remain constants when $\kappa=\omega(1)$. Hence, it would be intriguing to achieve a constant competitive ratio in the regime $m=\omega(n)$. Moreover, relaxing the assumptions that servers outnumber requests and all requests must be served gives rise to an interesting variant where request rejection is permitted—–a setting that bears similarities to the multi-unit prophet inequality problem [JMZ22]. Finally, it is appealing to relax the triangle inequality satisfied by the cost function; in Appendix D, we make progress toward this direction by generalizing some of our results to the case where the cost function only approximately satisfies the triangle inequality.